Grade 12Inorganic Chemistry

the equivalent weigt of a metal is 30 . what volume of chloride would combine at 27 C and 750 mm pressure with 0.6 gm of the metal ? (aqueous tension at 27 C = 26.6mm)

Mohammed Patel
13 Years agoGrade 12

1 Answer

Akash Kumar Dutta
ApprovedApproved Tutor Answer13 Years ago

.6/30=x/22.4
get x
apply P1.v1/t1=p2v2/t2..where p2=750-26.6
v1=x at stp.
and get v2 ans